OpenStudy (anonymous):

our family of 6 went to a Italian restaurant for dinner. Everyone had a sub, either meatball for $9 or chicken parmesan for $11. If the total bill was $58, how many people ordered each type of sub?

OpenStudy (anonymous):

9x+11y=58 x+y=6 here x-->people having meatball and y-->people havin parmesan

OpenStudy (anonymous):

4 had meatball and 2 had chicken parmesan

OpenStudy (anonymous):

x+y=6 x=6-y 9x+11y=58 subtitution x=6-y 9(6-y)+11y=58 54-9y+11y=58 2y=58-54 y= 4/2 y=2 x+2=6 x= 6-2 x=4 4meatball , 2 chicken

OpenStudy (anonymous):

use Good Answer button:)